OpenStudy (anonymous):

If a shirt costs $39 and the sales tax is 5%, how much is the total cost of buying the shirt? A. $40.95 B. $39.05 C. $44.00 D. $41.85

OpenStudy (anonymous):

.05 * 39 = 1.95 1.95 + 39 = 40.95

OpenStudy (anonymous):

The taxes are 5% of 39$ : \[39$\times(5\div100) = 1,95$\] So the toal price is \[39$+1,95$ = 40,95$\]